What are the no. of electrons present in 1c of charge?
 

Grade:12

1 Answers

Arun
4 years ago

One coulomb is the fundamental SI unit of charge. It is defined as the charge in one Ampere of current passed for one second.

Now for an electron, the amount of charge it carries is 1.6*10^-19 C.

So for making a charge of 1 Coulomb, the number of electrons required will be:

Q = n*e

Q = 1 C and e = 1.6*10^-19

n = 1/(1.6*10^-19)

n = 6.25*10^18 number of electrons